### v1

# Performant near-term quantum combinatorial optimization

Abstract
We present a variational quantum algorithm for solving combinatorial optimization problems with linear-depth circuits. Our algorithm uses an ansatz composed of Hamiltonian generators designed to control each term in the target combinatorial function, along with parameter updates following a modified version of quantum imaginary time evolution. We evaluate this ansatz in numerical simulations that target solutions to the MAXCUT problem. The state evolution is shown to closely mimic imaginary time evolution, and its optimal-solution convergence is further improved using adaptive transformations of the classical Hamiltonian spectrum, while resources are minimized by pruning optimized gates that are close to the identity. With these innovations, the algorithm consistently converges to optimal solutions, with interesting highly-entangled dynamics along the way. This performant and resource-minimal approach is a promising candidate for potential quantum computational advantages on near-term quantum computing hardware.
Introduction – Combinatorial optimization is widely viewed as a scientific computing challenge where quantum computers may make a significant impact Preskill (2018); Herman et al. (2023a); Abbas et al. (2023). Research on this topic has been motivated by several encouraging results obtained for the quantum approximate optimization algorithm (QAOA) Farhi et al. (2014); Hogg (2000) as well as its variants Hadfield et al. (2019); Herman et al. (2023b); Ushijima-Mwesigwa et al. (2021); Tate et al. (2023); Bärtschi and Eidenbenz (2020); Wurtz and Love (2021a). These include theoretical analyses Wurtz and Love (2021b); Díez-Valle et al. (2023); Wurtz and Love (2022), which have shown high QAOA approximation ratios for the SK model Farhi et al. (2022) and for Maximum-cut on large girth graphs Basso et al. (2021), as well as a variety of encouraging results on hardware Ebadi et al. (2022); Maciejewski et al. (2023); Harrigan et al. (2021); Pagano et al. (2020); Niroula et al. (2022); Shaydulin and Pistoia (2023); Pelofske et al. (2023) and numerical simulations Zhou et al. (2020a); Lotshaw et al. (2021); Herrman et al. (2021); Golden et al. (2023); Crooks (2018), such as a numerical scaling advantage for the low-autocorrelation binary sequence (LABS) problem Shaydulin et al. (2023). Modified QAOA-like ansätze have also determined ways to improve performance by introducing additional parameterized gates Herrman et al. (2022), including gates with counter-diabatic Hamiltonians Chandarana et al. (2022a) or based on Hamiltonians selected to maximize the energy convergence rate Zhu et al. (2022a), as well as many others Blekos et al. (2023).
Despite encouraging progress there are several drawbacks to current QAOA-based approaches. The primary limitation is that deep circuits are required for high performance Farhi et al. (2022); Lykov et al. (2023); Lotshaw et al. (2023); Marwaha (2021) while noise at these depths places severe limitations on theoretical and experimental performance Stilck França and Garcia-Patron (2021); Lotshaw et al. (2022); Weidenfeller et al. (2022); De Palma et al. (2023). Graph instances can even be designed so that any given QAOA depth $p$ cannot outperform classical algorithms in the noiseless case Bravyi et al. (2020). One reason large depths are required is that a single QAOA layer only entangles qubits that are adjacent in a given problem graph; hence the algorithm can require significant numbers of layers to build correlations across all qubits Farhi et al. (2020a, b), and further layers to remove this entanglement and approach the optimal classical solution Dupont et al. (2022). It has been argued that entanglement actually serves as a barrier to QAOA performance due to the prohibitive depths required to remove it Chen et al. (2022). This is concerning—entanglement should be helping our quantum algorithms, not hurting them. There have been approaches to heuristically modify the QAOA ansatz to reduce the depth or improve performance, such as Refs. Herrman et al. (2022); Chandarana et al. (2022a); Zhu et al. (2022a); Blekos et al. (2023). While these and related approaches do give noteworthy empirical improvements, they often do not improve on the theoretical $p\to\infty$ guarantee of standard QAOA, and can also introduce their own problems in terms of more involved parameter optimizations, as well as losing QAOA’s physical motivation in terms of an approximation to an optimal annealing algorithm Brady et al. (2021a, b). It is desirable to find other physically motivated approaches that do not have these shortcomings.
In this letter, we present a new algorithm for quantum combinatorial optimization that overcomes several issues of previous approaches such as QAOA. Building on recent work on counter-diabatic inspired ansätze Chandarana et al. (2023); Xu et al. (2024), we show that our algorithm is guaranteed to be able to prepare the optimal solution at linear depth given suitable parameters. We use parameters that are optimized through a systematic procedure, based on variational quantum imaginary time evolution and with an energy-dependent nonlinear objective that isolates the optimal solutions during training. Numerical simulations show that the final states at different training iterations transition from approximate product states, to volume-law entangled states, and to final states confined to the optimal solution subspace. Our ansatz accesses these highly entangled states during training of short-depth circuits, rather than in deep circuits, as in QAOA. Furthermore, we find that a small number of iterations succeed in identifying optimal solutions to a wide variety of problems we consider. The high-performance, low-depth, and modest training costs distinguish our algorithm as a prime candidate for successful near-term quantum optimization.
Combinatorial optimization – The goal of a combinatorial optimization problem is to minimize a cost function $C(\bm{z})$ with an $N$-bit argument $\bm{z}=(z_{1},\ldots,z_{N})$. These problems can often be formulated as quadratic unconstrained binary optimization (QUBO) with $C(\bm{z})=\sum_{i<j}J_{i,j}z_{i}z_{j}+\sum_{i}h_{i}z_{i}$. For quantum computing these problems are mapped to a diagonal Hamiltonian Lucas (2014)
$H_{c}=\sum_{\alpha}P_{\alpha},\ \ \ H_{c}\left|\bm{z}\right\rangle=C(\bm{z})\left|\bm{z}\right\rangle$ (1)
where $P_{\alpha}$ are Pauli-$Z$ strings and the optimal solution is the ground state of $H_{c}$. Here we focus on the standard benchmarking problem of weighted MAXCUT. Given a graph $G=(V,E)$ with edge weights $\{w_{i,j}\}$, the goal is to bipartition the vertices such that the summed weight of edges with endpoints in different sets is maximized. The optimal bipartition $\bm{z}_{\mathrm{opt}}$ minimizes $C(\bm{z})=\sum_{(i,j)\in E}w_{ij}z_{i}z_{j}$ with $z_{i}\in\{1,-1\}$ (physics convention, $H_{c}=\sum_{(i,j)\in E}w_{ij}Z_{i}Z_{j}$) or $C^{\prime}(\bm{z})=-\sum_{(i,j)\in E}w_{ij}(z_{i}+z_{j}-2z_{i}z_{j})/2$ with $z_{i}\in\{0,1\}$ (computer science convention, $H_{c}=-\sum_{(i,j)\in E}w_{ij}(1-Z_{i}Z_{j})/2$). Performance is quantified by the approximation ratio
$\mathrm{AR}=\frac{\langle C\rangle}{C_{\mathrm{opt}}}$ (2)
where $C_{\mathrm{opt}}$ is the optimal cost value. The different conventions are offset by a constant $\sum_{(i,j)\in E}w_{ij}/2$ which affects the approximation ratio. Following previous work, we use the physics convention for the Sherrington-Kirkpatric model, and the computer science convention for other cases.
Variational Ansatz – A recent series of papers by Xi Chen and collaborators has considered QAOA-like ansatze that incorporate counterdiabatic terms to improve convergence Chandarana et al. (2022a); Xu et al. (2024), recently culminating in a highly-parameterized ansatz for protein folding problems Chandarana et al. (2023) that we build upon here. The main ingredients of said ansatz are $ZY$ rotations, which directly rotate between the between the ±$\pm 1$ sector of $ZZ$ terms of the Hamiltonian. For MAXCUT, their prescription Xu et al. (2024) kept only $ZY$ rotations corresponding to $ZZ$ terms that also occur in $H_{c}$. We found that optimization of that ansatz commonly struggled to converge, so we introduce a fully connected ansatz of the form
$\left|\Psi(\theta)\right\rangle=\prod_{j=\{(r,q):r<q\}\in|V|}\exp(i\theta_{j}Z_{r}Y_{q})\left|+\right\rangle^{\otimes}$ (3)
where $j$ indexes all unique pairs of vertices. The ordering of the vertices is chosen by summed vertex weight as explained in S1. For problems such as MAXCUT consisting of only quadratic $ZZ$ terms, the ground state sector will consist of one or more states locally equivalent (by $X$ flips) to the $N$-qubit GHZ state. It can be shown that (3) can reach any such GHZ pair up to inconsequential relative phases, and is overdetermined for such a mapping. Each $ZY$ rotation is implementable on quantum hardware in terms of either two CNOTs or one two-qubit gate of the form $\exp(-i\theta ZZ)$ as found on certain trapped ion architectures such as Quantinuum devices. The two-qubit gate count for this ansatz is meager: Defining $N_{p}=N(N-1)/2$, we need only $2N_{p}$ CNOT gates, or half that if using arbitrary $ZZ$ rotations. In comparison, a single layer of QAOA requires a parameterized two-qubit rotation for each $ZZ$ term in $H_{c}$, which is less expensive than our ansatze for the least connected graphs and equivalent for fully connected graphs; we will show the cost of our ansatz can be minimized in later sections.
Parameter Optimization – System agnostic, classical parameter optimizers are known to suffer from common pitfalls like local minima Anschuetz and Kiani (2022) and vanishing gradients McClean et al. (2018). To improve performance we turn to a novel implementation of quantum imaginary time (QITE). Prior implementations of QITE relied on McLachlan’s principle McArdle et al. (2019, 2020); Alam et al. (2023), minimizing
$\bigg{|}\bigg{|}\left(\frac{\partial}{\partial\tau}+H_{c}-E_{\tau}\right)|\Psi(\vec{\theta})\rangle\bigg{|}\bigg{|}$ (4)
with respect to angle variations. Variationally evolving in imaginary time on a quantum computer by Eq. 4 requires $N_{p}(N_{p}+1)/2$ circuits per iteration, with most relying on controlled ansatz preparation McArdle et al. (2019, 2020).
Figure 1: Variational imaginary time evolution, following (3) and (6), closely matches exact imaginary time evolution for an example 8 vertex, fully connected graph with edge weights sampled from U(0,1) and timestep $\Delta\tau=0.1$.
We instead derive an alternative of QITE that requires fewer function evaluations and no controlled ansatz preparation. We notice that it is simpler instead to enforce that the expectation value of each $P_{\alpha}$ appearing in $H_{c}$ should evolve according the imaginary time as,
$\frac{\partial\langle P_{\alpha}\rangle}{\partial\tau}$ $=\sum_{j}2\mathfrak{R}\left(\langle\Psi(\vec{\theta})|P_{\alpha}\frac{\partial|\Psi(\vec{\theta})\rangle}{\partial\theta_{j}}\right)\dot{\theta_{j}}$
$=-\langle\Psi(\vec{\theta})|\{P_{\alpha},H_{c}-E_{\tau}\}|\Psi(\vec{\theta})\rangle\,.$ (5)
Performing these evaluations for each $P_{\alpha}$ in $H_{c}$, we arrive at the matrix equation,
$G\cdot\dot{\vec{\theta}}=D$ (6)
where
$G_{\alpha,j}=\mathfrak{R}\left(\langle\Psi(\vec{\theta})|P_{\alpha}\frac{\partial|\Psi(\vec{\theta})\rangle}{\partial\theta_{j}}\right)$ (7)
$D_{\alpha}=-\frac{1}{2}\langle\Psi(\vec{\theta})|\{P_{\alpha},H_{c}-E_{\tau}\}|\Psi(\vec{\theta})\rangle\,.$ (8)
For Ising like Hamiltonians encountered in $H_{c}$, all $P_{\alpha}$ commute, so $D$ can be evaluated in a single circuit. Further each column of $G$, corresponding to a given parameter derivative, requires only two circuits. To see this note that variations of $P_{\alpha}$ with respect to a single Pauli rotation [from Eq. (3)] have the form
$\langle P_{\alpha}\rangle(\theta_{j}+\delta)=a_{0}+b_{0}\cos(2\delta)+2G_{\alpha,j}\sin(2\delta)\,.$ (9)
Thus to determine $G$ in QUBO, we need only $2N_{p}$ circuit evaluations plus one circuit for energy. Then equation (6) can be inverted numerically to determine $\dot{\vec{\theta}}$ and parameters can be updated with a simple forward Euler step $\vec{\theta}\rightarrow\vec{\theta}+\Delta\tau\dot{\vec{\theta}}$. Fig. 1 demonstrates the success of our variational imaginary time (VAR-IT) approach for an 8-vertex fully connected MAXCUT problem with uniform random weights sampled from U(0,1) and the ansatz Eq. (3).
Improving Convergence – For large $N$ graphs, using $H_{c}$ directly generally imposes harsh requirements on step size as different parameters have rates of change that differ by orders of magnitude. Additionally, for many families of MAXCUT systems of interest, the spectral gap becomes vanishingly small with large $N$, leading to slow convergence in imaginary-time-based methods. To mitigate these difficulties, we leverage the manifestly diagonal form of $H_{c}$, introducing the sigmoidal transformation
$f(H_{c})=\left(1+\exp\left(\frac{-\sigma_{0}(H_{c}-E_{\tau})}{4\sigma_{\tau}^{2}}\right)\right)^{-1}$ (10)
where $E_{\tau}$, $\sigma_{\tau}$, and $\sigma_{0}$ are the current energy, current standard deviation, and initial standard deviation in the $\left|+\right\rangle^{\otimes N}$ state. This function retains the same optimal cut, but stretches the spectrum of states within a standard deviation of the current energy, while pushing farther states towards energies of zero and one. As the imaginary time evolution decreases $E_{\tau}$ and $\sigma_{\tau}$, even dense regions of excited states nearly degenerate with the ground state will approach a gap of 1 in $f(H_{c})$. In the rest of work, Eq. (8) is evaluated using $f(H_{c})$ and $\langle f(H_{c})\rangle$, instead of $H_{c}$ and $E_{\tau}$, respectively.
Figure 2: Average error convergence for 100 random instances of (a) 3 regular, (b) NWS, and (c) SK graphs with 16 vertices. The orange line is VAR-IT, the green line is (3) optimized with ADAM, and the purple line shows the optimized error from 8-layer QAOA [from Dupont and Sundar (2024); Dupont ].
Numerical and finite shot considerations makes exactly evaluating derivatives in Eq. (6) a poorly posed task, which manifests most strongly at the end of the imaginary time evolution. We therefor proceed by solving (6) [with $f(H_{c})$ in place of $H_{c}$] until the lowest energy cut measured achieves half the measurements, discarding the effect of singular values $\lambda_{\nu}<0.01\lambda_{max}$ in $G$, where $\lambda_{max}$ is the largest singular value. Once this is achieved, we perform Jacobi updates based on a single parameter while holding all the others constant, according to Parrish et al. (2019); performing this update once for all parameters brings the energy and norm on optimal solutions numerically to one, apart from a single Sherrington-Kirkpatric (SK) instance where VAR-IT converged to a near-degenerate excited state.
Numerical Results – We explore three prototypical problem types, 3 regular graphs, random Newman-Watts-Strogatz (NWS) small-world graphs, and the SK spin glass model with more details in Supplemental Information. These represent linear connectivity, non-trivial clustering type connectivity, and fully connected problem instances, respectively. In Fig. 2 we show the approximation ratio error produced by our VAR-IT algorithm and by Eq. (3) using the ADAM optimizer, averaged over one-hundred 16-vertex instances of the three problem types. VAR-IT always converges to the optimal energy, unlike the ADAM optimization. In Supplemental Information we show the fraction of instances that obtain $<1\%$ norm on the optimal energy after 100 iterations is increasing rapidly with size for the ADAM optimization of NWS and SK graphs; it is clear that ADAM will suffer a decreasing, and likely vanishing, probability of obtaining the optimal energy as the system size increases, whereas there is no signature of similar failure in VAR-IT in the sizes explored here. Additionally, VAR-IT converges in significantly fewer iterations than ADAM. Also plotted are average approximation ratios for $p=8$ QAOA, from Ref. Dupont and Sundar (2024); Dupont , which compare unfavorably to both ADAM and VAR-IT. All of the results in Fig. 2 can be improved by the “relaxed-rounding” postprocessing method of Ref. Dupont and Sundar (2024), but still VAR-IT outperforms these leading alternatives, as shown in Supplemental Information.
Figure 3: Excluding rotations with $\theta_{j}<\epsilon$. a) maximum ansatz gate counts in terms of equivalent layers of QAOA, and b) iterations required to reach an approximation ratio of 1.
When the number of $ZZ$ terms in $H_{c}$ scales linearly in problem size, QAOA requires $\sim N$ rotation gates while our ansatz requires $\sim N^{2}$ gates. To combat this shortcoming we explore omitting $ZY$ rotations if angles are smaller than a threshold $\epsilon$ at any point in the optimization. In Fig. 3, we apply this adaptive procedure to 100 3-regular graphs for $\epsilon=0.05,0.10,0.15$. The maximum number of rotations exploited during each graphs solution in terms of equivalent QAOA layers, is averaged and plotted in the left panel. For 3-regular graphs and a given $\epsilon$, the circuit cost appears to scale linearly at 2-3 layers of QAOA, with a minor increase in the number of iterations until convergence relative to the full ansatz. In Supplemental Information, iterations required to converge scale linearly in problem size for SK and NWS, similarly as seen for 3 Regular in Fig. 3(b).
Entanglement – If our algorithm produces large amounts of entanglement then it suggests the algorithm cannot be simulated at scale and therefor has potential for a quantum advantage, while low amounts of entanglement would seem to preclude quantum advantage due to simulability by tensor networks or related methods Zhou et al. (2020b).
We analyzed the entanglement entropy $S$ produced by our ansatz across 100 instances of the Sherrington-Kirkpatrick model at various $N$. For each instance we chose a random bipartition into sets of $N/2$ qubits and computed the entanglement entropy of the bipartition for the state produced at each step in the training, with averages over the instances shown in Fig. 4(a) and colored bands denoting the standard error of the mean. There is a transition from states with low entanglement at small numbers of iterations, where states are close to the initial product state, to highly entangled states midway through training, and finally to states with decreasing entropy as the training approaches optimal solutions; the curves terminate at the average iteration at which the states reach fidelity $F=1/2$ with the optimal solution subspace, at which point the algorithm switches to Jacobi iterations to isolate the optimal solutions. Fig. 4(b) shows the maximum entanglement entropy during these training iterations scales as a volume law $S\approx aN+b$. The behavior is similar to what is observed in QAOA Dupont et al. (2022); Chen et al. (2022), except with the crucial difference that we see this transition while training a shallow circuit, rather than in the dynamics of a deep circuit. Harnessing this entanglement at low depth distinguishes our algorithm as a more practical approach to solving combinatorial problems with a potential advantage.
Figure 4: (a) Entanglement entropy $S$ during training of the SK model. (b) Average maximum entropy fit by $S(N)=aN+b$, with $a=0.181\pm 0.004$ and $b=0.35\pm 0.05$; the vertical axis is the same as (a).
Conclusion – We developed an algorithm for quantum combinatorial optimization that makes several important advances over leading alternatives, in terms of optimization procedure and iterations, convergence behaviors, and circuit depth; widespread success of the algorithm is evident in simulations of diverse problems with up to 20 qubits. Key features of our approach include counterdiabatic Hamiltonian generators capable of preparing optimal solutions, training through QITE with a nonlinear objective that emphasizes the ground state, and resource-efficient generation of volume-law entanglement during training. We expect the low depth and minimal training requirements of our algorithm to enable quantum computers to solve combinatorial optimization problems with several tens of qubits in the immediate future, with extensions to larger sizes and scaling analyses beyond classical simulability as quantum computers continue to develop. The low depth, minimal resource requirements, high entanglement, and superior empirical performance make our approach a leading contender for near-term quantum computational advantage in combinatorial optimization.
Acknowledgements – We thank Joseph Wang, Eugene Dumitrescu, Jim Ostrowski, Ryan Bennink, George Siopsis, Rizwanul Alum, Noah Bauer, and Rebekah Herrman for discussions of variational quantum optimization algorithms. We thank Maxime Dupont for discussion and sharing numerical results from Ref. Dupont and Sundar (2024). T.D.M was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center and devised the ansatz, optimization algorithm, performed the simulations, and contributed to writing. P.C.L. was supported by the DARPA ONISQ program under award W911NF-20-20051 and contributed to design, analysis, writing, and entanglement computations. This research used resources of the Compute and Data Environment for Science (CADES) at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.